You want a seat on the board of directors of Four Keys, Inc. The company has 225,000 shares of stock outstanding and the stock sells for $66 per share. There are currently 3 seats up for election. If the company uses cumulative voting, how many shares do you need to guarantee that you will be elected to the board?
Answer #1

Shares Needed = [No. of shares outstanding / (N + 1)] + 1; N =  No. of seats open for election

= [225,000 / (3 + 1)] + 1

= 56,250 + 1 = 56,251 shares